OpenStudy (anonymous):

HELP AWARDING METAL!! Find all Solutions to the Equation. sin2x + sin x = 0

OpenStudy (bibby):

\(\sin^2x + \sin x = 0 \) is that it?

OpenStudy (anonymous):

Yes

OpenStudy (bibby):

you can factor it like \(\sin x(\sin x + 1)=0\) when is a product = to 0?

OpenStudy (anonymous):

Is that One solution?

OpenStudy (anonymous):

thats it :)